OpenStudy (bootifulwowzer):

Please help me! (x/x-2)+(x-1/x+1)=-1

OpenStudy (supercalifragalisticexspeaalli):

what do u need me to do

OpenStudy (bootifulwowzer):

I do not know how to solve this problem. Is it even solvable? I need to find x

OpenStudy (supercalifragalisticexspeaalli):

um yea i can sole it for u

OpenStudy (bootifulwowzer):

I need the steps too

OpenStudy (supercalifragalisticexspeaalli):

@Hero

OpenStudy (bootifulwowzer):

I am confused

OpenStudy (coopergang):

Solutions, X=0, x=1 alternate form 2x^2-2x+1)/(x-2)(x+1)=-1

OpenStudy (bootifulwowzer):

I do not know how to get that answer however^

OpenStudy (coopergang):

so what

OpenStudy (bootifulwowzer):

I need the steps leading up to the answer as well as the solution itself

OpenStudy (latinc):

you have to isolate x or create a quadratic equation like ax^2+bx+c = 0

OpenStudy (bootifulwowzer):

Okay I have that: x^2-3x+1=0

OpenStudy (bootifulwowzer):

What comes next?

OpenStudy (latinc):

now factor the equation like (x+a)(x+b)

OpenStudy (latinc):

or

OpenStudy (bootifulwowzer):

It cannot be simplified though

OpenStudy (latinc):

 use this equation